A linear supply curve is QXS = 5 - 4Px. Does it follow the law of supply? Why?
Answer #1

Law of supply states that as the price of a good increases ,the quantity of goods supplied will also increase and vice versa ,ceteris paribus.

If Q^S= 5-4Px ,then dQ^S/dPx= - 4 ,which simply means that an increase in the price level by 1 dollar decreases quantity supplied by 4 units . So,there is a negative relationship between price and quantity supplied.

If you have any doubt,feel free to ask.

Don't forget to thumbs up if you like the explanation.